OpenStudy (anonymous):

what is the sum of no between 100 and 1000 which are divisible by 14

OpenStudy (anonymous):

\[ 14*8=112\] \[14*71=994\] therefore there are \[64 (71-7)\] numbers between 100 and 1000 which are divisible by 14 since all these numbers are in arithmetic progression \[\sum=64/2*(112+994)\] \[= 35392\]